# Single-shot and measurement-based quantum error correction via fault complexes

### v1

@@ Abstract
A $(w,c)-$overlapping window decoder determines in each round a correction for a window of $w\in\mathbb{N}^{+}$ rounds and commits a correction to $c\leq w$ rounds (see Refs. [48, 50] for a more detailed description). Here, one round constitutes a check node and a bit node layer; see Fig. 1. For the 3D toric code, the effective distance of the decoding window then becomes $\min(wL,L^{2})$, see  Appendix B. However, $wL$ is the weight of time-like logical errors and in this section we consider memory experiments, i.e., only space-like logical errors cause logical failures (see the End Matter for stability experiment simulations). State-of-the-art results in decoding higher-dimensional topological codes [22] employ a single-stage decoding approach that is recovered using a $(1,1)$-overlapping window decoder, see Appendix D. We find that increasing $w$ to $2$ or $3$ significantly increases the sustainable threshold of 3D and 4D toric codes compared to $w=1$ when using belief propagation (BP) plus ordered statistics decoding (OSD) [51, 52]; see Appendix D for details concerning the simulations. For $w=3$, the thresholds for phenomenological Pauli noise of approximately $9.65\%$ (3D) and $5.9\%$ (4D) surpass all previous results [53, 54, 55, 56, 57, 58, 21, 22], approaching the thresholds achieved with the optimal window choice $w=L$, see Fig. 2(a) and Fig. 2(c). We note that for the 3D toric code, we only consider primal faults, as the dual side of the fault complex does not have the single-shot property.
Additionally, we investigated the threshold under a noise model for Xanadu’s photonic architecture based on Gottesman-Kitaev-Preskill (GKP) qubits [59, 1, 2], with cluster state construction and error model details provided in [4]. For the 4D toric code, we observe a threshold of approximately $10.35$ dB, comparable to that of the 2D variant [1, 2]. The 3D code achieves a threshold of approximately $7.95$ dB. In both cases, the $w=3$ decoder archives thresholds comparable to the $w=L$ decoder, see Fig. 2(b) and Fig. 2(d).
@@
While larger decoding regions increase decoding time, recent advances provide solutions. Findings on localized statistics decoding [60], a parallelizable OSD variant, and on belief propagation for higher-dimensional toric codes [61] suggest time-efficient decoding with minimal performance reduction, supporting the practical applicability of our results.
@@
Conclusion.— We introduced fault complexes, which represent dynamic QEC protocols and provide a convenient characterization of foliated codes. Through the explicit calculation of the homology groups of the fault complex, we gained insights into stability experiments and lattice surgery, and we derived improved decoding protocols for higher-dimensional topological codes such as the 4D toric code. The increased threshold values and asymptotic space-time advantage warrant serious consideration of the practical implementation of the 4D code in potentially compatible architectures such as photonics [1, 2, 3], trapped ions [62] or neutral atoms [63]. Algorithmic fault tolerance [64] offers asymptotically lower overhead, but it demands complex, correlated decoding. In contrast, the 4D toric code is more amenable to windowed decoding, and while the formal analysis requires decoding over $O(\sqrt{d})$ rounds, our numerical findings indicate that fewer layers might suffice in practice (see also [47]). On an intermediate time scale, the lower threshold requirements of 3D toric codes make them a promising experimental candidate to validate noise models and expected performance.
@@
On a broader scale, it is worth exploring what MBQC can teach us about CBQC, e.g., by extending our formalism to circuit-level noise. Reformulating Li’s [32] LDPC representation of fault-tolerant quantum circuits is potentially straightforward in our approach, as many of their results in [32, Sec. VIII] are already structurally equivalent to Eq. 8. It would also be interesting to explore the extension of our formalism to subsystem codes [65, 66, 67], especially those with single-shot error correction [36, 45, 68]. Additionally, symplectic chain complexes may be explored to further incorporate non-CSS codes. See Appendix E for some ideas on these extensions.
We applied our formalism to foliation, i.e., constructing fault complexes using tensor products of chain complexes. In future work, we plan to investigate alternative product constructions of fault complexes using, e.g., the balanced product [43] and the lifted product [69]. Given that these products are known to yield CSS codes with optimal parameters [42], they may also yield improved fault complexes. In addition, by using alternative linear codes in the fault complex construction, we may hope to algebraically recover FTGS beyond foliation [7, 29].
@@ Acknowledgements.
Acknowledgments.— The authors would like to thank Ben Brown, Christopher Chubb, Austin Daniel, Arthur Pesah, Alex Townsend-Teague, and Dominic Williamson for insightful discussions. We thank Rafael Alexander, J. Eli Bourassa, and Eric Sabo for feedback on an earlier draft of this manuscript. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Colleges and Universities. Computations were performed on the Niagara supercomputer at the SciNet HPC Consortium. SciNet is funded by Innovation, Science and Economic Development Canada; the Digital Research Alliance of Canada; the Ontario Research Fund: Research Excellence; and the University of Toronto.
Appendix A End Matter
Stability Experiments.— To further support our claim that $3D$ and $4D$ toric codes are compatible with single-shot lattice surgery, we perform simulations of stability experiments. Stability experiments can be used as a proxy for simulating lattice surgery because they probe the accuracy of reliably inferring the outcomes of an extensive number of stabilizers [23]. We perform numerical simulations over $32$ decoding rounds for a 3D toric code of linear size $L$ (with periodic boundaries in time to avoid finite-size effects). In this case, the corresponding fault complex is equivalent to the chain complex that describes an asymmetric 4D toric code with dimensions $L\times L\times L\times w$ for each decoding window. As described in the main text and in Appendix D, decoding proceeds with an $(w,1)$ overlapping window decoder. In this case, the effective primal fault distance is $d_{1}=wL$, i.e., for $w\leq L$ the distance is limited by the size of the decoding window. But even for $w=1$ we expect to observe an error threshold based on the distance scaling $d\sim L$. The $w=1$ case corresponds to single-shot lattice surgery, but we note that any $w=O(1)$ corresponds to constant time overhead lattice surgery. Our results are summarized in Fig. 3. As expected from the distance calculation, we observe an error threshold for all values of $w$, with the logical error rate decreasing with increasing $L$ below the threshold. By increasing the decoding window size $w$ from $w=1$ in Fig. 3(a), to larger, but constant, values of $w=2$ in Fig. 3(b) and $w=3$ in Fig. 3(c) we observe increased error thresholds in the stability experiments. We note that choosing $w$ adaptively on the code size, e.g., $w=L$ balances time-like and space-like errors. In practice, this allows one to find a trade-off between the number of syndrome measurement rounds required for lattice surgery, and thus the time overhead for logical gates, and the achievable logical error rates of the gate. We emphasize that these observations are in stark contrast to stability experiments with non-single shot codes, as performed by Gidney [23] on the 2D surface code, where the effective distance is independent of $L$. For two codes with $L_{2}>L_{1}$, the code with $L_{1}$ will always have a lower error rate than the code with $L_{2}$ if decoded over the same window size $w$.
Figure 3: Logical error rate for the time-like logical correlation of a 3D toric code for different window sizes of the overlapping window decoder. In (a) $w=1$, (b) $w=2$, and (c) $w=3$. Vertical black dashed lines indicate the fitted threshold. The shading indicates a confidence interval, indicating hypotheses whose likelihoods are within a factor of 1000 of the maximum likelihood estimate.
Appendix B Derivation of logical correlations and errors
Here we derive the expressions for logical errors/correlations and fault distances given in the main text and discuss their interpretation in more detail. We consider the fault complex =×$\mathcal{F}=\mathcal{R}\times\mathcal{C}$, where $\mathcal{C}$ describes a CSS code and $\mathcal{R}$ describes the repetition code whose parity-check matrix has a redundant row [^1] One can think of this parity-check matrix as being the vertex-to-edge adjacency matrix of the cycle graph., as the main text considers a special case of this. Explicitly,
$R=\begin{bmatrix}1&1&0&\ldots&0&0\\ 0&1&1&\ldots&0&0\\ \vdots&&&\ddots&&\vdots\\ 0&0&0&\ldots&1&1\\ 1&0&0&\ldots&0&1\\ \end{bmatrix}.$ (13)
@@ Appendix B Derivation of logical correlations and errors
$H_{j}(\mathcal{F})\cong H_{0}(\mathcal{R})\otimes H_{j}(\mathcal{C})\oplus H_{1}(\mathcal{R})\otimes H_{j-1}(\mathcal{C}).$ (14)
@@
$H^{j}(\mathcal{F})\cong H^{0}(\mathcal{R})\otimes H^{j}(\mathcal{C})\oplus H^{1}(\mathcal{R})\otimes H^{j-1}(\mathcal{C}).$ (15)
@@
$\begin{split}H_{0}(\mathcal{R})&=R_{0}/\operatorname{im}R,\\ H_{1}(\mathcal{R})&=\ker R/\{\mathbf{0}\}\cong\ker R,\\ H^{0}(\mathcal{R})&=\ker R^{T}/\{\mathbf{0}\}\cong\ker R^{T},\\ H^{1}(\mathcal{R})&=R_{1}/\operatorname{im}R^{T},\end{split}$ (16)
@@
$\begin{split}&\operatorname{im}R=\operatorname{im}R^{T}\\ &=\operatorname{span}\{(1,1,0,\ldots,0),\ldots,(0,\ldots,0,1,1)\}\end{split}$ (17)
@@
$\begin{split}H_{0}(\mathcal{R})&\cong\{\mathbf{0},(1,0,\ldots,0)\},\\ H_{1}(\mathcal{R})&\cong\{\mathbf{0},\mathbf{1}\},\\ H^{0}(\mathcal{R})&\cong\{\mathbf{0},\mathbf{1}\},\\ H^{1}(\mathcal{R})&\cong\{\mathbf{0},(1,0,\ldots,0)\}.\end{split}$ (18)
Using Eqs. 14 and 18 we have
$\begin{split}H_{i}(\mathcal{F})&\cong H_{0}(\mathcal{R})\otimes H_{i}(\mathcal{C})\oplus H_{1}(\mathcal{R})\otimes H_{i-1}(\mathcal{C})\\ &=\langle((1,0,\ldots,0)\otimes\ell_{Z},\mathbf{0}),(\mathbf{0},\mathbf{1}\otimes h)\rangle,\end{split}$ (19)
@@
$\begin{split}H^{i}(\mathcal{F})&\cong H^{0}(\mathcal{R})\otimes H^{i}(\mathcal{C})\oplus H^{1}(\mathcal{R})\otimes H^{i-1}(\mathcal{C})\\ &=\langle(\mathbf{1}\otimes\ell_{X},\mathbf{0}),(\mathbf{0},(1,0,\ldots,0)\otimes g)\rangle,\end{split}$ (20)
@@
$\begin{split}H_{i+1}(\mathcal{F})&\cong H_{0}(\mathcal{R})\otimes H_{i+1}(\mathcal{C})\oplus H_{1}(\mathcal{R})\otimes H_{i}(\mathcal{C})\\ &=\langle((1,0,\ldots,0)\otimes g^{\prime},\mathbf{0}),(\mathbf{0},\mathbf{1}\otimes\ell_{Z}^{\prime})\rangle,\end{split}$ (21)
@@
$\begin{split}H^{i+1}(\mathcal{F})&\cong H^{0}(\mathcal{R})\otimes H^{i+1}(\mathcal{C})\oplus H^{1}(\mathcal{R})\otimes H^{i}(\mathcal{C})\\ &=\langle(\mathbf{1}\otimes h^{\prime},\mathbf{0}),(\mathbf{0},(1,0,\ldots,0)\otimes\ell_{X}^{\prime})\rangle,\end{split}$ (22)
@@
Recall that the number of primal and dual logical correlations (or errors) of a fault complex are defined to be $k_{i}=\dim H_{i}(\mathcal{F})$ and $k_{i+1}=H_{i+1}(\mathcal{F})$, respectively. Using Eq. (14), we obtain the general expression for these quantities
@@
which in our case further simplifies via Eq. (18) giving
$\begin{split}k_{i}&=\dim H_{i}(\mathcal{C})+\dim H_{i-1}(\mathcal{C}),\\ k_{i+1}&=\dim H_{i+1}(\mathcal{C})+\dim H_{i}(\mathcal{C}),\end{split}$ (23)
@@
Let us now consider the interpretation of the logical correlations and errors in terms of the CSS code described by $\mathcal{C}$. We concentrate on the primal case as the dual case is analogous. Note that each $\ell_{Z}$ describes the support of a logical $Z$ operator (of the CSS code) and each $\ell_{X}$ describes the support of a logical $X$ operator. Therefore, symplectic pairs of logical operators in the CSS code lift to symplectic pairs of logical correlations and errors in the fault complex. If $\partial_{i}$ is full rank, then $\operatorname{im}\partial_{i}=C_{i-1}$ and $\ker\partial_{i}^{T}=\{\mathbf{0}\}$. In this case, the only possible choices for $g$ and $h$ are the zero vector, and therefore the generators of the homology groups all derive from the logical operators of the CSS code. Now suppose that $C_{i}$ is rank deficient. In this case, $g$ describes a subset of $X$-type stabilizer generators whose product is the identity, but which is not an $X$-type meta-check of the code. And $h$ is a vector in $C_{i-1}$ with trivial meta-syndrome, but which is not the syndrome of a $Z$-type error (see the main text for an example). If the code does not have $X$-type meta-checks (i.e., $\mathcal{C}$ is length-2) then $g$ can be any subset of $X$-type stabilizer generators with trivial product and $h$ can be any vector in $C_{i-1}$ that is not the syndrome of a $Z$-type error.
@@
We note that one can always construct a metacheck matrix for a rank-deficient parity-check matrix. However, at least in topological codes, the metachecks are usually defined to be the relations comprising $O(1)$ stabilizer generators, i.e., the metacheck matrix has low-weight rows (and low-weight columns as it turns out). On the other hand, the spacelike invariants such as $g$ and $g^{\prime}$ above are relations comprising an extensive number of stabilizer generators. The rationale for this distinction is that if each stabilizer measurement is noisy, then combining an extensive number of measurement outcomes will yield no useful information in the limit of large code size. Indeed, this is what makes it non-trivial to preserve the value of these logical correlations in QEC protocols.
Let us also comment explicitly on the case of a single-shot CSS code with rank-deficient parity-check matrices that is completely described by a length-2 chain complex. Here, there is no explicit notion of metachecks. As noted above, metachecks can be constructed based on the linear dependencies in the parity-check matrices. The resulting metacheck matrices can be interpreted as classical codes with distances $d_{M_{X}}$ and $d_{M_{Z}}$. Our results therefore extend to codes where $d_{M_{X}}$ and $d_{M_{Z}}$ are non-trivial.
B.1 Fault distances
The fault distances of $\mathcal{F}$ can essentially be read off Eqs. 19 and 22, but we provide a derivation here for completeness. For an arbitrary length-1 chain complex $\mathcal{R}$, Zeng and Pryadko showed in Ref. [14] that the product complex =×$\mathcal{F}=\mathcal{R}\times\mathcal{C}$ has homological distances
$\begin{split}d_{j}(\mathcal{F})&=\min[d_{0}(\mathcal{R})d_{i}(\mathcal{C}),d_{1}(\mathcal{R})d_{i-1}(\mathcal{C})],\\ d_{j}(\mathcal{F}^{T})&=\min[d_{0}(\mathcal{R}^{T})d_{j}(\mathcal{C}^{T}),d_{1}(\mathcal{R}^{T})d_{j-1}(\mathcal{C}^{T})],\end{split}$ (24)
@@ B.1 Fault distances
$\begin{split}d_{0}(\mathcal{R})=d_{1}(\mathcal{R}^{T})=1,\\ d_{1}(\mathcal{R})=d_{0}(\mathcal{R}^{T})=\delta,\end{split}$ (25)
@@
$\begin{split}d_{i}&=\min[d_{i}(\mathcal{C}),\delta d_{i-1}(\mathcal{C})],\\ d_{i+1}&=\min[\delta d_{i+1}(\mathcal{C^{T}}),d_{i}(\mathcal{C^{T}})].\end{split}$ (26)
Appendix C Single-shot decoding for fault complexes
@@
We consider the fault complex =×$\mathcal{F}=\mathcal{R}\times\mathcal{C}$, where $\mathcal{R}$ describes a distance $\ell$ repetition code with the parity-check matrix $R$ and $\mathcal{C}$ is a length-4 chain complex describing a CSS code with meta-checks. For a CSS code of this type, we have
$\begin{split}H_{X}&=\partial_{2}^{C},\quad H_{Z}^{T}=\partial_{3}^{C},\\ M_{X}&=\partial_{1}^{C},\quad M_{Z}^{T}=\partial_{4}^{C},\end{split}$ (27)
@@
$\mathcal{C}=C_{5}\stackrel{{\scriptstyle\partial_{5}}}{{\rightarrow}}C_{4}\stackrel{{\scriptstyle\partial_{4}}}{{\rightarrow}}C_{3}\stackrel{{\scriptstyle\partial_{3}}}{{\rightarrow}}C_{2}\stackrel{{\scriptstyle\partial_{2}}}{{\rightarrow}}C_{1}\stackrel{{\scriptstyle\partial_{1}}}{{\rightarrow}}C_{0},$ (28)
@@
$\partial_{j}=\left(\begin{array}[]{c|c}\mathbbm{1}_{r}\otimes\partial^{C}_{j}&R\otimes\mathbbm{1}_{n_{j-1}}\\ \hline\cr 0&\mathbbm{1}_{c}\otimes\partial^{C}_{j-1}\end{array}\right).$ (31)

### v2

@@ Abstract
A $(w,c)$-overlapping window decoder determines in each round a correction for a window of $w\in\mathbb{N}^{+}$ rounds and commits a correction to $c\leq w$ rounds; see Refs. [54, 56] for a more detailed description. Here, one round constitutes a check node and a bit node layer; see Fig. 1. For the 3D toric code, the effective distance of the decoding window then becomes $\min(wL,L^{2})$ [37]. However, $wL$ is the weight of time-like logical errors, and in this section, we consider memory experiments, i.e., only space-like logical errors cause logical failures; see [37] for stability experiment simulations. State-of-the-art results in decoding higher-dimensional topological codes [24] employ a single-stage decoding approach that is recovered using a $(1,1)$-overlapping window decoder [37]. We find that increasing $w$ to $2$ or $3$ significantly increases the sustainable threshold of 3D and 4D toric codes compared to $w=1$ when using belief propagation (BP) plus ordered statistics decoding (OSD) [57, 58]; see [37] for details concerning the simulations. For $w=3$, the thresholds for phenomenological Pauli noise (see [37] for a photonic noise model) of approximately $9.65\%$ (3D) and $5.9\%$ (4D) surpass all previous results [59, 60, 61, 62, 63, 64, 23, 24, 65], approaching the thresholds achieved with the optimal window choice $w=L$; see Fig. 2(a) and Fig. 2(b). We note that for the 3D toric code, we only consider primal faults, as the dual side of the fault complex does not have the single-shot property.
@@
While larger decoding regions increase decoding time, recent advances provide solutions. Findings on localized statistics decoding [66], a parallelizable OSD variant, and belief propagation for higher-dimensional toric codes [65] suggest time-efficient decoding with minimal performance reduction, supporting the practical applicability of our results.
Figure 2: Sustainable thresholds for 3D (a) and 4D (b) toric codes under phenomenological Pauli noise. Markers represent the threshold for a $(w,1)$-overlapping window decoder as a function of noisy syndrome rounds. The black dashed line shows the threshold for the optimal window choice $w=L$; see [37] for details.
@@
Conclusion.— We introduced fault complexes to represent dynamic QEC protocols and characterize foliated codes. This approach provided insights into stability experiments and lattice surgery, and yielded improved decoding protocols for higher-dimensional topological codes like the 4D toric code. The increased thresholds and space-time overhead advantages make 4D toric codes promising for compatible architectures like photonics [1, 2, 3, 4, 5, 6], trapped ions [67], and neutral atoms [68]. While algorithmic fault tolerance [69] offers alower overhead, it requires complex decoding; 4D toric codes allow simpler windowed decoding. Meanwhile, the lower threshold requirements of 3D toric codes make them suitable for experimental validation in the near term.
@@
Our formalism can extend to circuit-level noise by, for example, reformulating Li’s [36] LDPC representation of fault-tolerant quantum circuits (many of their results in [36, Sec. VIII] are already structurally equivalent to Eq. 7). Future work could explore the fault complexes of subsystem codes [70, 71, 72] (especially those with single-shot error correction [42, 51, 73]), symplectic chain complexes for non-CSS codes, and alternative product constructions like balanced [49] and lifted [74] products. In addition, by using alternative linear codes in the construction, we may hope to algebraically recover FTGS beyond foliation [9, 33]. See [37] for some intial ideas on these extensions.
@@ Acknowledgements.
Acknowledgments.— The authors would like to thank Ben Brown, Christopher Chubb, Austin Daniel, Arthur Pesah, Tom Stace, Alex Townsend-Teague, and Dominic Williamson for insightful discussions. We thank Rafael Alexander, J. Eli Bourassa, and Eric Sabo for feedback on an earlier draft of this manuscript. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Colleges and Universities. Computations were performed on the Niagara supercomputer at the SciNet HPC Consortium. SciNet is funded by Innovation, Science and Economic Development Canada; the Digital Research Alliance of Canada; the Ontario Research Fund: Research Excellence; and the University of Toronto.
Supplementary Information for
“Single-shot and measurement-based quantum error correction via fault complexes”
S1 Derivation of logical correlations and errors
Here, we derive the expressions for logical errors/correlations and fault distances given in the main text and discuss their interpretation in more detail. We consider the fault complex $\mathcal{F}=\mathcal{R}\times\mathcal{C}$, where $\mathcal{C}$ describes a CSS code and $\mathcal{R}$ describes the repetition code whose parity-check matrix has a redundant row [^2] One can think of this parity-check matrix as being the vertex-to-edge adjacency matrix of the cycle graph., as the main text considers a special case of this. Explicitly,
$R=\begin{bmatrix}1&1&0&\ldots&0&0\\ 0&1&1&\ldots&0&0\\ \vdots&&&\ddots&&\vdots\\ 0&0&0&\ldots&1&1\\ 1&0&0&\ldots&0&1\\ \end{bmatrix}$. (S1)
@@
$H_{j}(\mathcal{F})\cong H_{0}(\mathcal{R})\otimes H_{j}(\mathcal{C})\oplus H_{1}(\mathcal{R})\otimes H_{j-1}(\mathcal{C}).$ (S2)
@@
$H^{j}(\mathcal{F})\cong H^{0}(\mathcal{R})\otimes H^{j}(\mathcal{C})\oplus H^{1}(\mathcal{R})\otimes H^{j-1}(\mathcal{C}).$ (S3)
@@
$\begin{split}H_{0}(\mathcal{R})&=R_{0}/\operatorname{im}R,\\ H_{1}(\mathcal{R})&=\ker R/\{\mathbf{0}\}\cong\ker R,\\ H^{0}(\mathcal{R})&=\ker R^{T}/\{\mathbf{0}\}\cong\ker R^{T},\\ H^{1}(\mathcal{R})&=R_{1}/\operatorname{im}R^{T},\end{split}$ (S4)
@@
$\begin{split}&\operatorname{im}R=\operatorname{im}R^{T}\\ &=\operatorname{span}\{(1,1,0,\ldots,0),\ldots,(0,\ldots,0,1,1)\}\end{split}$ (S5)
@@
$\begin{split}H_{0}(\mathcal{R})&\cong\{\mathbf{0},(1,0,\ldots,0)\},\\ H_{1}(\mathcal{R})&\cong\{\mathbf{0},\mathbf{1}\},\\ H^{0}(\mathcal{R})&\cong\{\mathbf{0},\mathbf{1}\},\\ H^{1}(\mathcal{R})&\cong\{\mathbf{0},(1,0,\ldots,0)\}.\end{split}$ (S6)
Using Eqs. S2 and S6 we have
$\begin{split}H_{i}(\mathcal{F})&\cong H_{0}(\mathcal{R})\otimes H_{i}(\mathcal{C})\oplus H_{1}(\mathcal{R})\otimes H_{i-1}(\mathcal{C})\\ &=\langle((1,0,\ldots,0)\otimes\ell_{Z},\mathbf{0}),(\mathbf{0},\mathbf{1}\otimes h)\rangle,\end{split}$ (S7)
@@
$\begin{split}H^{i}(\mathcal{F})&\cong H^{0}(\mathcal{R})\otimes H^{i}(\mathcal{C})\oplus H^{1}(\mathcal{R})\otimes H^{i-1}(\mathcal{C})\\ &=\langle(\mathbf{1}\otimes\ell_{X},\mathbf{0}),(\mathbf{0},(1,0,\ldots,0)\otimes g)\rangle,\end{split}$ (S8)
@@
$\begin{split}H_{i+1}(\mathcal{F})&\cong H_{0}(\mathcal{R})\otimes H_{i+1}(\mathcal{C})\oplus H_{1}(\mathcal{R})\otimes H_{i}(\mathcal{C})\\ &=\langle((1,0,\ldots,0)\otimes g^{\prime},\mathbf{0}),(\mathbf{0},\mathbf{1}\otimes\ell_{Z}^{\prime})\rangle,\end{split}$ (S9)
@@
$\begin{split}H^{i+1}(\mathcal{F})&\cong H^{0}(\mathcal{R})\otimes H^{i+1}(\mathcal{C})\oplus H^{1}(\mathcal{R})\otimes H^{i}(\mathcal{C})\\ &=\langle(\mathbf{1}\otimes h^{\prime},\mathbf{0}),(\mathbf{0},(1,0,\ldots,0)\otimes\ell_{X}^{\prime})\rangle,\end{split}$ (S10)
@@
Recall that the number of primal and dual logical correlations (or errors) of a fault complex are defined to be $k_{i}=\dim H_{i}(\mathcal{F})$ and $k_{i+1}=H_{i+1}(\mathcal{F})$, respectively. Using Eq. (S2), we obtain the general expression for these quantities
@@
which in our case further simplifies via Eq. (S6), giving
$\begin{split}k_{i}&=\dim H_{i}(\mathcal{C})+\dim H_{i-1}(\mathcal{C}),\\ k_{i+1}&=\dim H_{i+1}(\mathcal{C})+\dim H_{i}(\mathcal{C}),\end{split}$ (S11)
@@
Let us now consider the interpretation of the logical correlations and errors in terms of the CSS code described by $\mathcal{C}$. We concentrate on the primal case as the dual case is analogous. Note that each $\ell_{Z}$ describes the support of a logical $Z$ operator (of the CSS code) and each $\ell_{X}$ describes the support of a logical $X$ operator. Therefore, symplectic pairs of logical operators in the CSS code lift to symplectic pairs of logical correlations and errors in the fault complex. If $\partial_{i}$ is full rank, then $\operatorname{im}\partial_{i}=C_{i-1}$ and $\ker\partial_{i}^{T}=\{\mathbf{0}\}$. In this case, the only possible choices for $g$ and $h$ are the zero vector, and therefore the generators of the homology groups all derive from the logical operators of the CSS code. Now suppose that $C_{i}$ is rank-deficient. In this case, $g$ describes a subset of $X$-type stabilizer generators whose product is the identity, but which is not an $X$-type meta-check of the code. And $h$ is a vector in $C_{i-1}$ with trivial meta-syndrome, but which is not the syndrome of a $Z$-type error (see the main text for an example). If the code does not have $X$-type meta-checks (i.e., $\mathcal{C}$ is length-2) then $g$ can be any subset of $X$-type stabilizer generators with trivial product, and $h$ can be any vector in $C_{i-1}$ that is not the syndrome of a $Z$-type error.
@@
We note that one can always construct a metacheck matrix for a rank-deficient parity-check matrix by Gaussian elimination. However, at least in topological codes, the metachecks are usually defined to be the relations comprising $O(1)$ stabilizer generators, i.e., the metacheck matrix has low-weight rows (and low-weight columns as it turns out). On the other hand, the spacelike invariants such as $g$ and $g^{\prime}$ above are relations comprising an extensive number of stabilizer generators. The rationale for this distinction is that if each stabilizer measurement is noisy, then combining an extensive number of measurement outcomes will yield no useful information in the limit of large code size. Indeed, this is what makes it non-trivial to preserve the value of these logical correlations in QEC protocols.
Let us also comment explicitly on the case of a single-shot CSS code with rank-deficient parity-check matrices that is completely described by a length-2 chain complex. Here, there is no explicit notion of metachecks. As noted above, metachecks can be constructed based on the linear dependencies in the parity-check matrices. The resulting metacheck matrices can be interpreted as classical codes with distances $d_{M_{X}}$ and $d_{M_{Z}}$. Our results therefore extend to codes where $d_{M_{X}}$ and $d_{M_{Z}}$ are non-trivial. However, such procedures generally lead to higher-weight rows. Determining meta-checks of minimal weight is expected to be (at least) NP-hard, in analogy to determining the minimum distance of a classical code.
S1.1 Fault distances
The fault distances of $\mathcal{F}$ can essentially be read off Eqs. S7 and S10, but we provide a derivation here for completeness. For an arbitrary length-1 chain complex $\mathcal{R}$, Zeng and Pryadko showed in Ref. [20] that the product complex $\mathcal{F}=\mathcal{R}\times\mathcal{C}$ has homological distances
$\begin{split}d_{j}(\mathcal{F})&=\min[d_{0}(\mathcal{R})d_{i}(\mathcal{C}),d_{1}(\mathcal{R})d_{i-1}(\mathcal{C})],\\ d_{j}(\mathcal{F}^{T})&=\min[d_{0}(\mathcal{R}^{T})d_{j}(\mathcal{C}^{T}),d_{1}(\mathcal{R}^{T})d_{j-1}(\mathcal{C}^{T})],\end{split}$ (S12)
@@
$\begin{split}d_{0}(\mathcal{R})=d_{1}(\mathcal{R}^{T})=1,\\ d_{1}(\mathcal{R})=d_{0}(\mathcal{R}^{T})=\delta,\end{split}$ (S13)
@@
$\begin{split}d_{i}&=\min[d_{i}(\mathcal{C}),\delta d_{i-1}(\mathcal{C})],\\ d_{i+1}&=\min[\delta d_{i+1}(\mathcal{C^{T}}),d_{i}(\mathcal{C^{T}})].\end{split}$ (S14)
S2 Single-shot decoding for fault complexes
@@
We consider the fault complex $\mathcal{F}=\mathcal{R}\times\mathcal{C}$, where $\mathcal{R}$ describes a distance $\ell$ repetition code with the parity-check matrix $R$ and $\mathcal{C}$ is a length-4 chain complex describing a CSS code with meta-checks. For a CSS code of this type, we have
$\begin{split}H_{X}&=\partial_{2}^{C},\quad H_{Z}^{T}=\partial_{3}^{C},\\ M_{X}&=\partial_{1}^{C},\quad M_{Z}^{T}=\partial_{4}^{C},\end{split}$ (S15)
@@
$\mathcal{F}=F_{5}\stackrel{{\scriptstyle\partial_{5}}}{{\rightarrow}}F_{4}\stackrel{{\scriptstyle\partial_{4}}}{{\rightarrow}}F_{3}\stackrel{{\scriptstyle\partial_{3}}}{{\rightarrow}}F_{2}\stackrel{{\scriptstyle\partial_{2}}}{{\rightarrow}}F_{1}\stackrel{{\scriptstyle\partial_{1}}}{{\rightarrow}}F_{0},$ (S16)
@@
$\partial_{j}=\left(\begin{array}[]{c|c}\mathbbm{1}_{r}\otimes\partial^{C}_{j}&R\otimes\mathbbm{1}_{n_{j-1}}\\ \hline\cr 0&\mathbbm{1}_{c}\otimes\partial^{C}_{j-1}\end{array}\right).$ (S19)